2p - q = 8
1/3p + 1/4q = 3 not sure how to make the fractions to whole number

Answer by stanbon(75887)

Multiply both sides of the 1st equation by "2" to get:
4p-2q=16
Multiply both sides of the 2nd equation by "12" to get:
4p+3q=36
Subtract one from the other to get:
5q=20
q=4
Substitute into the original equation to get:
2p-4=8
p=6
Cheers,
stan H.